1) Calculate the height of the rocket. It accelerates up at 10 m/s^2 for 3 seconds and then accelerates for -10 m/s^2 for 3 seconds until it reaches its peak.

2) A plane is doing a loop de loop. Where is G-force the strongest?
Answer #1

here,

for acceleration period ,

a1 = 10 m/s^2

t1 = 3 s

the height of rocket , h1 = 0 + 0.5 * a1 * t1^2

h1 = 0.5 * 10 * 3^2 m = 45 m

the final velocity , v1 = 0 + 10 * 3 = 30 m/s

for deacceleration period ,

the height covered , h2 = v1^2 /2a2

h2 = 30^2 /(2 * 3) = 150 m

the height of rocket , h = h1 + h2

h = 195 m